A solution is labeled 0.670 M HCN. What are the following in the solution? Express your answers to three significant figures even when only two are valid. Acid-base table [HCN]- [CN -] pH-
Answer #1

ans)

from acid base table

Ka for HCN =  4.0e-10

         HCN    +    H2O       H3O+     +     CN-
i       .670 0                 0
Δ        -x                               +x                +x
f       .670-x                           x                  x

4.0e-10 = x2 / (0.670-x)

2.68e-10 - 4.0e-10x = x2

1.64e-5 - 2.0e-5x = x

1.64e-5 = 1.00002x

1.64e-5 M = x

x = [H3O+] = [CN-] =1.64e-5M
pH = - log 1.64e-5 = 4.78

*[HCN] =0.670 M